A toy manufacturer uses 48,000 rubber wheels per year for his popular dump truck series. The firm makes its own wheels which it can produce at a rate of 800 per day. The toy trucks are assembled uniformly over the entire year. Carrying cost is $1 per wheel a year. Setup costs for a production run of wheels is $45. The firm operates 240 days per year. , Determine a) the optimum run size and b) minimum total annual cost for carrying and setup.
Answer #1

Answer:

Given data,

Annual demand (D) = 48000 wheels

Setup cost(S) = $45

Holding cost(H) = $1 per wheel per year

Daily production rate (p) = 800 per day

Daily demand rate (d) = Annual demand / Number of days in a year = 48000/240 = 200

a) Optimal run size (Q) = Sqrt of {2DS / H[1-(d/p)]}

= Sqrt of {(2 x 48000 x 45) / 1[1-(200/800)]}

= Sqrt of (4320000 / 0.75)

= Sqrt of 5760000

2400 wheels

b) Imax = (Q/p)(p-d) = (2400/800)(800-200) = 3 x 600 = 1800 wheels

Total annual cost for carrying and setup = Carrying cost + Setup cost

= [(Imax / 2)H] + [(D/Q)S]

= [(1800/2)1] + [(48000/2400)45]

= $900 + $900

$1800


answered by: Subrahmanyam golla